The width of rectangle is 1 ft less than the length. The area is 6ft^2. find the lenght and the width.
width is 2 ft
the length 3 ft.

Found 2 solutions by dyakobovitch, checkley77:
Answer by dyakobovitch(40)

Width=L-1
Length=L
Area=LW=L(L-1)=6

L^2-L-6=0
(L-3)(L+2)=0

So, L=3 [since -2 is an extraneous solution]
By substitution, Width=L-1=3-1=2

--
David

Answer by checkley77(12569)

W=L-1
WL=6
(L-1)L=6
L^2-L=6
L^2-L-6=0
(L-3)(L+2)=0
L-3=0
L=3 IS THE LENGTH OF THE RECTANGLE.
W=3-1=2 IS THE WIDTH OF THE RECTANGLE.
PROOF:
2*3=6
6=6